# Proper Name Diacritization for Arabic Wikipedia: A Benchmark Dataset

### v2

@@ 5.3 Manual Diacritization
The manual diacritization and quality checks were carried out by a native speaker of Arabic who is a trained linguist and highly experienced annotator. The annotation process initially was done in tandem with the finalization of the guideline with a team of the authors working jointly to optimize the quality of the annotation. The annotator was provided an Arabic word, along with its English gloss, and a proposed diacritization from GPT-4o after being refined by the automatic post-process described above. The annotations were carried on Google Sheets in a very simple setup. The annotator reviewed the proposed diacritization making changes where needed in accordance to the guidelines. The annotator made changes to 909 proposed lemmas out of 3,362 ($\sim$27%). In 213 instances (6.3% of all entries), there was a change connected with lemmatization: 74% relative involved the Al determiner, 22.5% a change in Alif-Hamza spelling, and 3.3% involving the demonym plural ending.
@@ 6.1 Experimental Setup
We perform computational experiments to perform the task of diacritization of proper nouns. For this, we prompt GPT-4o on all of the annotated dataset described in Section 5. We prompt the model with different input formats to assess its capabilities while giving it different levels of information: the inputs and the number of examples shown to the model (shots).
@@
The model is given a detailed description of the task to be performed. Our main experiments reflect all the information given to our annotator, where we provide the model with both the Arabic Input and the English Gloss (Arabic Input and Gloss). Additionally, we also experiment with a more constrained setup where the model is provided solely with the the Arabic Input (Arabic Input Only).
@@
In addition to the different inputs, we also consider further experiments where we supply the model with varying number of examples to learn from. Hence, in addition to just providing the input to diacritize (Zero-Shot), we also supply the model with a single example (One-Shot), and 80 examples (Few-Shot). The examples are randomly sampled from the AraPropComm data. However, since AraPropComm has fully lemmatized Arabic Inputs, we manually manipulated some of the examples to have a representation of clitic removal and Hamza normalization. Refer to Appendix A for a more detailed description of the prompts used. As a post-processing step, the outputs were ran through the same processing pipeline mentioned in Section 5.2. To evaluate the performance of the different experiments, we compute accuracy by measuring the exact match between the post-processed output and the gold-standard diacritization.
@@ 6.2 Results
Table 7 shows the results with different prompts.
Input Format Shots Accuracy
Arabic Input and Gloss Zero 46.49%
Arabic Input and Gloss One 61.90%
Arabic Input and Gloss Few 72.96%
Arabic Input Only Zero 36.67%
Arabic Input Only One 49.73%
Arabic Input Only Few 55.86%
Table 7: GPT-4o model Results measured by exact match accuracy.
@@
The results demonstrate that while diacritizing proper nouns remains a challenging task, incorporating the English gloss offers a valuable signal for the model. Notably, the best performance is achieved with few-shot prompting, showing the effectiveness of providing a diverse and representative sample.
@@ 6.3 Error Analysis
We analyzed errors from a sample of 1,010 output entries from the best performing setup from Section 6.2, and classified errors into several categories based on observed patterns. There were 740 (73.3%) exact matches (correct generations). Of the 270 (26.7%) errors, there were 175 cases where the error was only diacritization differences. See examples in Table 8. Upon further analysis of this class of errors, we found that the model overpredicts fathas (+25%) and shaddas (+96%), while underpredicting kasras (-18%) and sukuns (-23%), indicating imbalanced vowel modeling and overuse of gemination. The next largest class of errors, 60 cases, where those with spelling changes limited to the set of long vowel (and glides) letters \<ا> A, \<و> w, and \<ي> y (AWY). As we see in the examples in Table 8, the model has the tendency of dropping such letters rather than adding them. Another class of errors, 10 cases, are those with specific letter replacements such as \<ج> j $\leftrightarrow$ \<غ> $\gamma$, \<خ> x $\leftrightarrow$ \<ه> h, and \<ـة> $\hbar$leftrightarrow \<ـه> h. The final class of errors, 25 cases, are those with multiple changes happening at once. While these cases don’t match the gold reference, they are plausible and acceptable alternatives in most cases, especially in the context of linguistic variation discussed in Section 2. For example, the generated diacritization for \<بلاجيفيتش> blAjyfytš ‘Blažević’ as seen in Table 8, follows the common phenomena of breaking word initial complex onsets found in many spoken dialects of Arabic and in MSA. Another example is the entry \<ايكوميديا> AykwmydyA ‘Eco-Médias’, where the input follows a pronunciation-based transliteration while the generated form adhered to the orthography of the gloss. These observations indicate the need for exploring different modeling techniques that accommodate this special case of Arabic diacritization which in turn requires additional annotated examples.
@@ 7 Conclusion and Future Work
We presented a new 3,362-entry dataset of Arabic Wikipedia proper names annotated with gold-standard lemma diacritizations, paired with their English equivalents. This resource enables the joint study of diacritization and transliteration in a realistic setting characterized by ambiguity and spelling variation. We benchmarked GPT-4o on this task, providing insights into its capabilities and limitations. While the model performs reasonably well, especially on frequent names, it struggles with rarer entries and variant mappings. Looking ahead, we plan to expand the dataset with more diverse names, integrate it into a morphological analyzer, and explore fine-tuned models for diacritizing proper nouns in broader contexts. We also plan to fine-tune dedicated models for this task and develop more robust approaches to name ambiguity, especially with multiple valid diacritizations. We hope this resource advances Arabic NLP and name normalization in multilingual settings like Wikipedia.
@@
All data used in this project were sourced from publicly available Arabic Wikipedia entries and their corresponding English titles, in accordance with Wikimedia’s terms of use. The annotation process was conducted transparently and ethically, with fair compensation provided to the annotator. We make both the corpus and the annotation guidelines publicly accessible under an open license, supporting reproducibility and community collaboration. Our goal is to contribute a valuable resource for Arabic language processing and to aid the broader Wikimedia effort by enhancing the quality of Arabic Wikipedia entries. Finally, we acknowledge that all NLP tools and resources can be used with malicious intent; this is not our intention, and we categorically discourage it.
@@ Appendix A GPT-4o Prompts
In the system role, we provide the task description, and optionally, the few-shot demonstrations, when they are used. For the user role, we always provide the single instance to be diacritized. Table 9 lists all of the prompts used for the different settings.
@@
Table 10 shows a sample of the few-shot examples. These are formatted as a markdown table in the prompts.

### v3

@@ 5.3 Manual Diacritization
The manual diacritization and quality checks were carried out by a native speaker of Arabic from Egypt who is a trained linguist and a highly experienced annotator. The annotation process initially was done in tandem with the finalization of the guidelines with a team of the authors working jointly to optimize the quality of the annotation. The annotator was provided an Arabic word, along with its English gloss, and a proposed diacritization from GPT-4o after being refined by the automatic post-process described above. The annotations were carried on Google Sheets in a very simple setup. The annotator reviewed the proposed diacritization making changes where needed in accordance to the guidelines. The annotator made changes to 909 proposed lemmas out of 3,362 ($\sim$27%). In 213 instances (6.3% of all entries), there was a change connected with lemmatization: 74% relative involved the Al determiner, 22.5% a change in Alif-Hamza spelling, and 3.3% involving the demonym plural ending.
5.4 Inter-annotator Agreement
To assess the quality of our annotation and the consistency of our guidelines, we conducted an inter-annotator agreement study. A second annotator, a native Arabic speaker from Egypt, independently re-annotated a subset of 500 randomly selected samples from the dataset, utilizing the same annotation process and adhering to the same guidelines as the first annotator. Out of the 500 samples, the annotators fully agreed on 462 instances and disagreed on 38, resulting in an inter-annotator agreement rate of 92.4%. Table 7 presents the various types of inter-annotator disagreements along with their corresponding frequencies. Each row in the table represents a type of disagreement where the annotators selected different diacritics for the same word. For example, the first row shows instances where either one of the annotators chose a Kasra while the other selected a Sukun.
@@ 6.1 Experimental Setup
We perform computational experiments to perform the task of diacritization of proper nouns. For this, we prompt GPT-4o on all of the annotated dataset described in Section 5. We prompt the model with different input formats to assess its capabilities while giving it different levels of information: the inputs and the number of examples shown to the model (shots). We used default settings for optional parameters (e.g., temperature, top_p) from the gpt-4o-2024-11-20 snapshot. [^4] https://platform.openai.com/docs/api-reference/chat/create
@@
The model is given a detailed description of the task to be performed. Our main experiments reflect all the information given to our annotator, where we provide the model with both the Arabic Input and the English Gloss (Arabic + Gloss). Additionally, we also experiment with a more constrained setup where the model is provided solely with the the Arabic Input (Arabic Only).
@@
In addition to the different inputs, we also consider further experiments where we supply the model with varying number of examples to learn fromLABEL:footCP Hence, in addition to just providing the input to diacritize (Zero-Shot), we also supply the model with a single example (One-Shot), and 80 examples (Few-Shot). The examples are randomly sampled from the CP-Sama data. The one-shot and few-shot examples were selected once and reused across all model prompts. However, since CP-Sama has fully lemmatized Arabic Inputs, we manually manipulated some of the examples to have a representation of clitic removal and Hamza normalization. Refer to Appendix A for a more detailed description of the prompts used.
Post-processing
As a post-processing step, the outputs were ran through the same processing pipeline mentioned in Section 5.2. To evaluate the performance of the different experiments, we computed two metrics: accuracy by measuring the exact match between the post-processed output and the gold-standard diacritization and Levenshtein edit distance (Levenshtein, 1966) between the output and gold-standard diacritization.
@@ 6.2 Results
Input Format Shots Accuracy Distance
Arabic + Gloss Zero 46.5% 1.02
Arabic + Gloss One 61.9% 0.64
Arabic + Gloss Few 73.0% 0.41
Arabic Only Zero 36.7% 1.29
Arabic Only One 49.7% 0.86
Arabic Only Few 55.9% 0.71
Table 8: GPT-4o model results on CP-Wiki-D3K in terms of exact match accuracy and Levenshtein edit distance.
@@
The results demonstrate that while diacritizing proper nouns remains a challenging task, incorporating the English gloss offers a valuable signal for the model. Notably, the best performance is achieved with few-shot, showing the effectiveness of providing a diverse and representative sample. Table 8 shows the results with different prompts.
6.3 Interplay of Frequency, Similarity, and Accuracy
We investigated how lexical frequency and phonological similarity Freeman et al. (2006) affect model performance under our best configuration: few-shot prompting with Arabic + Gloss. The Freeman similarity score averaged a high 91% across the dataset, consistent with the transliteration focus of the task. We binned the data into 10 intervals based on Freeman score. The lowest-similarity bins (up to 50%), comprising only 3% of the data, contained mostly high-frequency named entities and translations, e.g., \<مصر> mSr for ‘Egypt’ and \<عملاق > $\varsigma$mlAq for ‘jötnar’. Despite their low similarity, this group achieved 13.9% higher accuracy and had, on average, 10 times the frequency compared to the rest of the data. The bins up to 90% similarity comprised 35% of the data; their average frequency is only 5% higher than the last bin, but their average accuracy is lower by 3.6% absolute. We found strong negative correlation between accuracy and edit distance (-0.95), confirming that higher accuracy aligns with fewer character edits. Frequency and Freeman score showed a moderate negative correlation (-0.69), likely due to high-frequency translated names. Freeman similarity and accuracy were also moderately negatively correlated (-0.70), indicating that frequent but phonetically dissimilar words are still predicted accurately. We analyzed performance across frequency quartiles (Q1 to Q4). Accuracy rose steadily from 65% in Q1 to 80% in Q4. The correlation between average frequency and accuracy across quartiles was 0.68, confirming the positive impact of frequency on model performance. Full analysis tables are presented in Appendix B.
@@ 6.4 Error Analysis
We analyzed errors from a randomly selected sample of 1,010 output entries from the best performing setup from Section 6.2, and classified errors into several categories based on observed patterns. There were 740 (73.3%) exact matches (correct generations). Of the 270 (26.7%) errors, there were 175 cases where the error was only diacritization differences. See examples in Table 9. Upon further analysis of this class of errors, we found that the model overpredicts Fathas (+25%) and Shaddas (+96%), while underpredicting Kasras (-18%) and Sukuns (-23%), indicating imbalanced vowel modeling and overuse of gemination. The next largest class of errors, 60 cases, were those with spelling changes limited to the set of long vowel (and glides) letters \<ا> A, \<و> w, and \<ي> y (AWY). As we see in the examples in Table 9, the model has the tendency of dropping such letters rather than adding them. Another class of errors, 10 cases, are those with specific letter replacements such as \<ج> j $\leftrightarrow$ \<غ> $\gamma$, \<خ> x $\leftrightarrow$ \<ه> h, and \<ة> $\hbar$leftrightarrow \<ه> h. The final class of errors, 25 cases, are those with multiple changes happening at once. While these cases don’t match the gold reference, they are plausible and acceptable alternatives in most cases, especially in the context of linguistic variation discussed in Section 2. For example, the generated diacritization for \<بلاجيفيتش> blAjyfytš ‘Blažević’ as seen in Table 9 (row 5), follows the common phenomena of breaking word initial complex onsets in many spoken dialects of Arabic and in MSA. Another example is the entry \<ايكوميديا> AykwmydyA ‘Eco-Médias’, where the input follows a pronunciation-based transliteration while the generated form adhered to the orthography of the gloss. These variations highlight the need for modeling techniques and evaluation metrics that account for this aspect of Arabic proper noun diacritization, which in turn requires additional annotated data.
@@ 7 Conclusion and Future Work
We presented a new 3,362 entry dataset of Arabic Wikipedia proper nouns annotated with gold-standard lemma diacritizations, paired with their English equivalents. This resource enables the joint study of diacritization and transliteration in a realistic setting characterized by ambiguity and spelling variation. We benchmarked GPT-4o on this task, providing insights into its capabilities and limitations. While the model performs reasonably well, especially on frequent names, it struggles with rarer entries and variant mappings. Looking ahead, we plan to expand the dataset with more diverse names, integrate it into a morphological analyzer, and explore fine-tuned models for diacritizing proper nouns in broader contexts. We also plan to fine-tune dedicated models for this task and develop more robust approaches to name ambiguity, especially with multiple valid diacritizations. We hope this resource advances Arabic NLP and name normalization in multilingual settings like Wikipedia.
@@
All data used in this project were sourced from publicly available Arabic Wikipedia entries and their corresponding English titles, in accordance with Wikimedia’s terms of use. The annotation process was conducted transparently and ethically, with fair compensation provided to the annotators. We make both the corpus and the annotation guidelines publicly accessible under an open license, supporting reproducibility and community collaboration. Our goal is to contribute a valuable resource for Arabic language processing and to aid the broader Wikimedia effort by enhancing the quality of Arabic Wikipedia entries. Finally, we acknowledge that all NLP tools and resources can be used with malicious intent; this is not our intention, and we categorically discourage it.
@@
Acknowledgments
We would like to express our sincere gratitude to Hamdy Mubarak for his valuable insights and his generous willingness to answer our questions throughout the course of this work. We thank Djellel Difallah for advice on initial data collection. We would like to also thank Bashar Alhafni and Mostafa Saeed for their insightful discussions and helpful conversations.
@@ Appendix A GPT-4o Prompts
In the system role, we provide the task description, and optionally, the few-shot demonstrations, when they are used. For the user role, we always provide the single instance to be diacritized. Table 10 lists all of the prompts used for the different settings. Table 11 shows a sample of the few-shot examples. These are formatted as a markdown table in the prompts.
@@
Appendix B Supplementary Interplay of Frequency, Similarity, and Accuracy
Freeman Bin Instances Instance % Frequency Matches Accuracy Distance
10% 6 0.2% 2,280,059 5 83.3% 0.17
20% 7 0.2% 454,346 6 85.7% 0.29
30% 23 0.7% 303,728 20 87.0% 0.22
40% 27 0.8% 690,729 23 85.2% 0.26
50% 26 0.8% 64,814 23 88.5% 0.12
60% 71 2.1% 30,274 45 63.4% 0.69
70% 164 4.9% 57,361 124 75.6% 0.37
80% 271 8.1% 22,803 185 68.3% 0.46
90% 587 17.5% 22,909 404 68.8% 0.52
100% 2,180 64.8% 60,343 1,619 74.3% 0.38
10–90% 1,182 35.2% 63,761 835 70.6% 0.48
10–50% 89 2.6% 496,420 77 86.5% 0.20
60–100% 3,273 97.4% 49,719 2,377 72.6% 0.42
All 3,362 100.0% 61,544 2,454 73.0% 0.41
Table 12: Accuracy, average frequency, and edit distance across Freeman similarity score bins.
Frequency Range Instances Average Freq. Matches Accuracy Avg. Freeman
Q1 (lowest 25%) 787 2 510 64.8% 91.1%
Q2 (25–50%) 893 25 627 70.2% 90.4%
Q3 (50–75%) 840 567 646 76.9% 91.2%
Q4 (highest 25%) 842 245,145 671 79.7% 89.7%
All 3,362 61,544 2,454 72.99% 90.6%
Table 13: Accuracy, Average Frequency, and average Freeman similarity scores across word frequency quartiles.